OpenStudy (anonymous):

Ms. Adams received a bonus check for $12,000. She decided to divide the money among three different investments. With some of the money, she purchased a municipal bond paying 5.8% simple interest. She invested twice the amount she paid for the municipal bond in a certificate of deposit paying 4.9% simple interest. Ms. Adams placed the balance of the money in a money market account paying 3.7% simple interest. If Ms. Adams' total interest for one year was $534, how much was placed in each account?

OpenStudy (anonymous):

A. municipal bond: $ 1500 certificate of deposit: $ 3000 money market: $ 7500 B. municipal bond: $ 2500 certificate of deposit: $ 5000 money market: $ 4500 C. municipal bond: $ 2000 certificate of deposit: $ 4000 money market: $ 6000 D. municipal bond: $ 1750 certificate of deposit: $ 3500 money market: $ 6750

OpenStudy (anonymous):

You throw a ball straight up from a rooftop. The ball misses the rooftop on its way down and eventually strikes the ground. A mathematical model can be used to describe the relationship for the ball's height above the ground, y, after x seconds. Consider the following data: x, seconds after ball is thrown y, ball's height, in feet, above the ground 1 114 2 146 4 114 Find the quadratic function y = ax2 +bx + c whose graph passes through the given points. A. y = -12x2 + 80x + 46 B. y = -10x2 + 60x + 64 C. y = -16x2 + 100x + 30 D. y = -16x2 + 80x + 50

OpenStudy (anonymous):

answer is d

OpenStudy (anonymous):

for the second one